### v1

# Coercivity Panorama of Dynamic Hysteresis

Abstract
We study the stochastic $\phi^{4}$ model under periodic driving by an external field $H$ at different scales of driving rate $v_{H}$, where the noise strength $\sigma$ quantifies the deviation of the system size from the thermodynamic limit. For large systems with small $\sigma$, we find the coercivity $H_{c}=H(\langle\phi\rangle=0)$ sequentially exhibits distinct behaviors with increasing $v_{H}$: $v_{H}$-scaling increase from zero, stable plateau ($v_{H}^{0}$), $v_{H}^{1/2}$-scaling increase, and abrupt decline to disappearance. The $H_{c}$-plateau reflects the competition between thermodynamic and quasi-static limits, namely, $\lim_{\sigma\to 0}\lim_{v_{H}\to 0}H_{c}=0$, and $\lim_{v_{H}\to 0}\lim_{\sigma\to 0}H_{c}=H^{*}$. Here, $H^{*}$ is exactly the first-order phase transition (FOPT) point. In the post-plateau slow-driving regime, $H_{c}-H^{*}$ scales with $v_{H}^{2/3}$. Moreover, we reveal a finite-size scaling for the coercivity plateau $H_{P}$ as $(H^{*}-H_{P})\sim\sigma^{4/3}$ by utilizing renormalization-group theory. These predicted scaling relations are demonstrated in magnetic hysteresis obtained with the Curie-Weiss model. Our work provides a panoramic view of the finite-time evolution of the stochastic $\phi^{4}$ model, bridges dynamics of FOPT and dynamic phase transition, and offers new insights into finite-time/finite-size effect interplay in non-equilibrium thermodynamics.
Figure 1: Hysteresis loops and coercivity panorama of the stochastic $\phi^{4}$ model. Hysteresis loops at (a): $v_{H}=10^{-5}$(solid), $4\times 10^{-4}$(dashed); (b): $v_{H}=0.01$(solid), $0.02$(dashed), $0.04$(dash-dotted); (c): $v_{H}=10$(solid), $20$(dashed), $40$(dash-dotted); (d): $v_{H}=100$(solid), $500$(dashed), $1500$(dash-dotted). (e) Coercivity $H_{c}$ as a function of driving rate $v_{H}$ with $\sigma=0.7$. (f) Coercivity landscape $H(v_{H},\sigma)$, where the noise strength from top to bottom is $\sigma=$ 0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0, 2.0. (g) A contour illustration of (f) after interpolation. In this plot, limited driving amplitude $H_{m}=50$ is applied in (a-e) while (f-g) not limited driving amplitude to ensure that the order parameter is always saturated; $a_{2}=-4$, and all quantities are nondimensionalized by $a_{4}$ and $\lambda$.
Introduction.—Hysteresis is ubiquitous across natural and engineered systems, spanning micro- to macro-scales. This phenomenon manifests in diverse fields, including magnetic and mechanical materials [1, 2, 3, 4, 5], electric and optical devices [6, 7], thermal systems [8, 9, 10], biological cycles [11], chemical reactions [12], and even economies [13]. Characterized by history-dependent dynamics, hysteresis reveals a system’s delayed response to external perturbations, where its current state relies on both instantaneous driving parameters and their temporal evolution. Within the realm of physics, magnetic hysteresis has garnered country-long attention due to its rich phenomenology, with extensive experimental and theoretical investigations [1, 14, 15, 16, 17]. It reveals fundamental magnetic properties—anisotropy, coercivity, and remanence—that underpin advancements in magnetic technologies, including permanent-magnet devices [18] and magnetic random access memories [19].
Notably, magnetic hysteresis loops exhibit diverse morphologies—some display steep characteristics indicative of first-order phase transitions (FOPT), while others feature smoother, thinner profiles [20, 21, 22]. A central focus in hysteresis research lies in how hysteresis loops evolve with driving parameters, for example, the rate-dependent deformations in Fig. 1(a-d). This issue encompasses not only the scaling laws of the hysteresis loop area with dynamic parameters [23, 24, 25, 26, 27, 28, 29] but also delves into phenomena like the dynamic phase transitions (DPT) [30, 31, 32]. Critical scaling analyses reveal a fundamental dichotomy: systems with and without quasi-static hysteresis exhibit distinct dynamical behaviors [23, 24], rooted in contrasting ergodic properties [33]. This disparity underscores the interplay between the system’s relaxation capability and driving dynamics [34, 35, 36, 37, 38, 39, 40, 41]. While mean-field theories predict a $2/3$ power law scaling for finite-time deviation from the quasi-static hysteresis in slow-driving FOPT systems [23, 42], experiments and finite-dimensional Ising models show a richer diversity of scaling behaviors [43, 44, 45, 46, 47, 48, 49, 50, 51].
Despite these significant advancements, current studies on hysteresis confront three critical challenges: i) In microscopic theories rooted in nonequilibrium statistical physics, scaling relations of dynamic hysteresis across diverse time scales remain largely unexplored [23, 52, 42, 53, 54]; ii) For phenomenological models in magnetism, such as the Jiles-Atherton model [55, 56, 57, 58], Preisach model [59, 60, 61, 62] and Stoner-Wohlfarth model [63, 64], there is a significant lack of understanding of the continuous transition between quasi-static and dynamic hysteresis [55, 65]. This raises a critical question: do hysteresis loops have a well-defined quasi-static limit, and how does it emerge with decreasing driving rate? iii) The comparison between experimentally measured and theoretically predicted scaling laws for hysteresis suffers from ambiguity [48, 66, 9, 10], primarily due to incomplete characterization of hysteresis dependence on driving rate across the entire accessible frequency regimes.
To address these gaps, we introduce the concept of coercivity panorama[Fig. 1(e)], derived from the driven stochastic $\phi^{4}$ model. This framework offers a unified lens to reconcile scaling relations across time scales and system sizes, highlighting competition between thermodynamic and quasi-static limits in interacting systems. Concomitantly, the derived order parameter’s evolution equation serves as a natural bridge between microscopic dynamics and macroscopic hysteretic behaviors, resolving long-standing theory-experiment mismatches in hysteresis modeling.
Coercivity plateau in periodically driven stochastic $\phi^{4}$ model.—For a system of interest described by the Landau model interacting with an external field $H$, its free energy density, as a function of the dimensionless order parameter $\phi$, reads [67]
$f_{4}(\phi,H)=\frac{1}{2}a_{2}\phi^{2}+\frac{1}{4}a_{4}\phi^{4}-H\phi,$ (1)
where the constants $a_{2}$ and $a_{4}$ encode the reduced temperature and intrinsic interaction strength, respectively. For $a_{2}<0$ at zero field ($H=0$), the free energy landscape exhibits $\mathbb{Z}_{2}$-symmetric bistability. Under weak applied fields satisfying $|H|<|H_{\mathrm{sp}}^{\pm}|$, this double-well potential becomes tilted while maintaining both local minima. Here, $H_{\mathrm{sp}}^{\pm}\equiv\pm H^{*}\equiv\pm\sqrt{(4/27)(-a_{2}^{3}/a_{4})}$ denotes the spinodal field, at which the metastable state ceases to exist and the metastable minimum merges with the unstable maximum at $\phi_{\mathrm{sp}}^{\pm}=\pm\phi^{*}\mp\text{sgn}(H_{\mathrm{sp}}^{\pm})\sqrt{-a_{2}/(3a_{4})}$. The superscripts $\pm$ denote transitions toward positive and negative $\phi$, respectively. The dynamics of $\phi=\phi(t)$ under a time-dependent field $H=H(t)$ can be described by the Langevin equation [68, 48]
$\frac{\partial\phi}{\partial t}=-\lambda\frac{\partial f_{4}(\phi,H)}{\partial\phi}+\zeta(t),$ (2)
where $\lambda$ sets a reference timescale and $\zeta(t)$ is a Gaussian white noise with amplitude $\sqrt{2\lambda\sigma^{2}}$ satisfying $\langle\zeta(t)\rangle=0$ and $\langle\zeta(t)\zeta(t^{\prime})\rangle=2\lambda\sigma^{2}\delta(t-t^{\prime})$, and $\sigma\geq 0$ denoting the noise strength. The system defined by Eqs. (1) and (2) is referred to as stochastic $\phi^{4}$ model throughout this work. In the deterministic limit ($\sigma\to 0$), the model describes a sharp FOPT [23, 53], exactly occurring at the spinodal fields $H=H_{\mathrm{sp}}^{\pm}$ under quasi-static driving between saturation fields of opposite signs. Finite driving rates delay the transition beyond $H=H_{\mathrm{sp}}^{\pm}$ due to dynamic hysteresis [23], while finite fluctuations ($\sigma>0$), representing deviations from the thermodynamic limit in our framework, cause premature escape from metastable states before reaching $H=H_{\mathrm{sp}}^{\pm}$. The competition between these finite-time and finite-size effects leads to complex nonequilibrium dynamics.
To directly capture the dynamics of the stochastic $\phi^{4}$ model, we first simulate its evolution with a discretized form of Eq. (2) (See Supplemental Materials [69] for details). Figures 1(a-d) display hysteresis loops for four different driving regimes, generated under a field swept triangularly between ±$\pm H_{m}$ at a rate $dH/dt=\lambda v_{H}$, where $H_{m}>0$ is the amplitude. Two key features emerge from these figures: (i) The hysteresis loop is symmetric with respect to the origin and the loop width increases monotonically with $v_{H}$ when the saturation field is smaller than $H_{m}$; however, at very fast driving rates, the loops shrink and form narrow cycles centered around the initial value of the order parameter. (ii) In the first two regimes, the hysteresis loops exhibit clear signatures of FOPT, characterized by a distinct turning point that triggers a rapid phase transition. In contrast, this feature becomes weaker and even vanishes in the latter two regimes.
In the field-driven transition from one stable phase to another, coercivity ($H_{c}$) is defined as the applied field value at which the ensemble-averaged order parameter crosses zero [70]. Figure 1(e) illustrates a coercivity panorama obtained from the periodically driven stochastic $\phi^{4}$ model, showing its evolution with increasing driving rate at a fixed field amplitude. This panorama strikingly reveals four distinct dynamic regimes: (i) a near-equilibrium regime where coercivity scales linearly with $v_{H}$; (ii) a plateau associated with FOPT; (iii) a post-plateau power-law scaling regime; and (iv) a DPT regime characterized by rapidly diminishing coercivity at high driving rates. Unlike the often-studied loop area response curves, which typically exhibit a long, decreasing tail after a maximum with increasing driving frequency [25, 26], this coercivity panorama uniquely focuses on the ascending regime, clearly identifying and explaining its distinct dynamic behaviors. Notably, the existence of the FOPT plateau is a novel finding presented in this study.
To show the FOPT plateau’s dependence on the system’s noise strength ($\sigma$), we plot the coercivity landscape $H_{c}(v_{H},\sigma)$ in Fig. 1(f), where $\sigma$ increases from top (deterministic limit $\sigma\to 0$) to bottom. We set the field amplitude to ensure the order parameter always saturates. As $\sigma$ decreases, the FOPT plateau extends toward the quasi-static limit ($v_{H}\to 0$), while increasing $\sigma$ narrows it until it vanishes, squeezed by adjacent regimes. This plateau arises from the competition between the lifetime of the metastable state [34] and the variance of the potential landscape. Intriguingly, a non-commutativity of limits emerges at the far left side of the FOPT plateau:
$\begin{cases}\lim_{\sigma\to 0}\lim_{v_{H}\to 0}H_{c}=0,\\ \lim_{v_{H}\to 0}\lim_{\sigma\to 0}H_{c}=H^{*},\end{cases}$ (3)
which is clearly illustrated by the contour plot in Fig. 1(g). Taking the quasi-static limit first implies that finite fluctuations are always present to ensure ergodicity. In contrast, taking the thermodynamic limit first eliminates fluctuations, causing the system to remain trapped in metastable states even under quasi-static driving. This distinction has also been reflected in the difference between deterministic mean-field results and fluctuating Monte Carlo simulations of the Ising model [45].
Finite-time scaling relations.—For the stochastic $\phi^{4}$ model under a linearly driven external field, the corresponding Fokker-Planck equation of Eq. (2) reads
$\frac{\partial P(\phi,t)}{\partial t}=\lambda\frac{\partial}{\partial t}\left[\frac{\partial f_{4}(\phi,H)}{\partial\phi}P(\phi,t)\right]+\lambda\sigma^{2}\frac{\partial^{2}P(\phi,t)}{\partial\phi^{2}},$ (4)
where $P(\phi,t)$ denotes the probability density of the system having the value $\phi$ at time $t$. Multiplying Eq. (4) by $\phi$ and integrating the resulting expression over all possible values of $\phi$, we obtain an ensemble-averaged equation,
$v_{H}\frac{d\langle\phi\rangle}{dH}=-a_{2}\langle\phi\rangle-a_{4}\langle\phi^{3}\rangle+H,$ (5)
where $dH=\lambda v_{H}dt$ has been used. This equation serves as the first main result of this Letter, capturing the evolution of measurable macroscopic quantities and linking microscopic stochastic models with phenomenological equations widely used in experimental magnetism [15].
By evaluating Eq. (5) at $\langle\phi\rangle=0$, we obtain the coercivity as
$H_{c}=v_{H}\evaluated{\frac{d\langle\phi\rangle}{dH}}_{\langle\phi\rangle=0}+a_{4}\evaluated{\langle\phi^{3}\rangle}_{\langle\phi\rangle=0}.$ (6)
Here, the first term on the right-hand side is related to the slope of the $\langle\phi\rangle-H$ curve at the coercivity point; the second term characterizes the skewness of the probability distribution of $\phi$ when $\langle\phi\rangle=0$. As a result of Eq. (6), three distinct scaling regimes of coercivity with respect to the driving rate are identified(See [69] for detailed derivations), as illustrated in Tab. 1.
Table 1: Finite-time coercivity scaling relations for the stochastic $\phi^{4}$ model.
near-equilibrium post-plateau fast-driving
$H_{c}\sim v_{H}$ $H_{c}-H_{P}\sim(v_{H}-v_{P})^{2/3}$ $H_{c}\sim v_{H}^{1/2}$
In the near-equilibrium regime, for non-divergent equilibrium value of $\evaluated{d\langle\phi\rangle/dH}_{\langle\phi\rangle=0}$, the scaling $H_{c}\sim v_{H}$ is denoted by the dashed line in Fig. 1(f). This linear response relation is universal for systems weakly driven out of equilibrium [71, 72, 73]. In the fast-driving regime, the dynamics is effectively governed by a passive relaxation process toward the mono-stable minimum induced by the continuously increasing field. The interplay between the relaxation timescale and the field variation rate leads to a $H_{c}\sim v_{H}^{1/2}$ scaling, which is denoted by the dash-dotted line in Fig. 1(f). While previous analytical studies on mean-field models often overlooked this fast-driving regime, this scaling relation finds support in existing numerical [74] and experimental evidence [75].
Finite-size scaling of the plateau.—Before analyzing the finite-time scaling behavior near the plateau, we first examine the plateau itself, clarifying its properties by investigating its dependence on system size and driving rate. Expanding the free energy (1) near the quasi-static spinodal points and retaining terms up to third order, the resulting discretized equation governing the variation of the order parameter reads
$\delta\varphi=-a_{3}\varphi^{2}v_{H}^{-1}\delta h+hv_{H}^{-1}\delta h+\sqrt{2\sigma^{2}v_{H}^{-1}\delta h}\;W,$ (7)
[…]
$H^{*}-H_{P}\sim\sigma^{4/3}\quad\text{at}\quad v_{P}\sim\sigma^{2},$ (8)
where $H_{P}$ denotes the height of the coercivity plateau, $H^{*}$ is that height in the deterministic limit, and $v_{P}$ is the reference driving rate at which the coercivity plateau is observed [77]. Up to now, the noise strength has been treated as a phenomenological parameter reflecting finite-size effects. The explicit dependence of $\sigma$ on system size $N$ can be illustrated through specific microscopic examples [47, 54].
Figure 2: Reduced coercivity after the FOPT plateau. (a) Coercivity reduced by the reference plateau coercivity value. Curves from right to left correspond to noise strengths $\sigma=$ 1.0, 0.8, 0.7, 0.5, 0.4, 0.3. (b) Coercivity reduced by the minimum observed coercivity value within the simulation timescale shown in Fig. 1(f). Curves from left to right correspond to $\sigma=$ 0.3, 0.2, 0.1, 0. Parameters: $a_{2}=-4$. All quantities are nondimensionalized by $a_{4}$ and $\lambda$.
Finite-size corrections to the $2/3$ power law scaling.—Unlike the universal scaling observed in the near-equilibrium and fast-driving regimes, the scaling behavior near the coercivity plateau is highly sensitive to system size and constrained by the finite timescales accessible in both simulations and experiments. By expanding the dynamic equation (2) around the coercivity plateau point $(v_{P},H_{P})$, we obtain a finite-time scaling as $H_{c}-H_{P}\sim(v_{H}-v_{P})^{2/3}$ [69], where both $v_{P}$ and $H_{P}$ depend on system sizes (8). In the deterministic limit $\sigma\to 0$, $H_{P}\rightarrow H^{*}$ and $v_{P}\rightarrow 0$, such that this scaling reduces to $H_{c}-H^{*}\sim v_{H}^{2/3}$. This is consistent with the scaling, $A-A_{0}\sim v_{H}^{2/3}$, characterized by subtracting the quasistatic hysteresis loop area $A_{0}$ from the finite-time area $A$ [23, 42, 54], because the loop geometry directly leads to $A-A_{0}\propto H_{c}-H^{*}$. This $2/3$-exponent scaling for finite $\sigma$ is demonstrated by the dotted lines in Fig. 2(a). However, for very small noise strengths, the genuine plateau point becomes unreachable within accessible simulation timescales [as shown in the upper set of Fig. 1(f)], making it challenging to clearly characterize the coercivity plateau. In such cases, it is natural to take the minimum observed coercivity $(H_{c}^{\mathrm{min}})$ as a quasi-static value [48]. Fig. 2(b) illustrates the scaling relation $H_{c}-H_{c}^{\mathrm{min}}$ with respect to $v_{H}$, which indicates a transient $2/3$-exponent scaling. This localized scaling behavior has also been reported in Ref. [54]. Depending on the accessible rate range and specific models, the prominent effective scaling exponent can be either smaller or larger than $2/3$ [48, 78]. This variability could explain observed deviations from deterministic mean-field predictions in the accessible slow-driving regime. In this sense, the global scaling landscape of coercivity provides a more robust and reliable characterization than the localized scaling behavior of the reduced coercivity near the FOPT plateau.
Magnetic coercivity in Curie-Weiss model.—To illustrate the finite-size scaling obtained from the stochastic $\phi^{4}$ model, we focus on the magnetic hysteresis of the Curie-Weiss model [79, 80, 54, 81], which is a solvable model of magnetism, also known as the mean-field Ising model (See End Matters for details). By comparing the roles of $\sigma$ and $N$ in the Fokker-Planck equations of the stochastic $\phi^{4}$ and Curie-Weiss models, one has $\sigma\sim N^{-1/2}$. By rescaling the axes in the $H_{c}-v_{H}$ plot according to the scaling relations in Eq. (8), it is found that the coercivity curves for different system sizes collapse onto a single universal curve, as shown in Fig. 3(a), confirming the validity of the finite-size scaling relations. This universal curve diverges at a critical driving rate that yields $H_{c}=H^{*}$. Here $H^{*}$ is the spinodal field of the Curie-Weiss model. In the infinite-size limit, this critical driving rate approaches zero, rendering the region to its left inaccessible. This demonstrates that the coercivity plateau resulting from taking the thermodynamic limit ($\sigma\rightarrow 0$) first in Eq.(3) does not vanish in the slow-driving regime. In contrast, for finite-size systems, coercivity values span both sides of this critical point, with the gently varying region at lower driving rates reflecting the presence of a coercivity plateau.
On the other hand, it is worth mentioning that the $v_{H}$-scaling in the near-equilibrium regime and the finite-size scaling relations associated with the FOPT plateau are universal across both the stochastic $\phi^{4}$ and Curie-Weiss models. However, in the fast-driving regime, despite sharing similar physical mechanisms, the specific scaling differs between the two models due to fundamental differences in their full dynamical equations. In the Curie-Weiss model, as shown in Fig. 3(b), the scaling behavior (yellow dash-dotted curve) in the fast-driving regime follows the Lambert W-function $W(v_{H})$ [69]. Moreover, between the FOPT plateau and the fast-driving regime, an intermediate scaling (dashed line), $H_{c}\sim v_{H}^{1/3}$ [69], emerges. While this scaling is absent as a sustained regime in the stochastic $\phi^{4}$ model, it is briefly passed through without being maintained over an observable timescale. This regime bridges the region governed by FOPT characteristics and the passive relaxation dynamics at high driving rates. These discrepancies highlight the sensitivity of far-from-equilibrium dynamics to model-specific details, even when the underlying physical picture remains similar. Consequently, different microscopic models may reveal more varied scaling relations in this regime  [52].
Figure 3: Coercivity of the Curie-Weiss model as a function of driving rate. (a) Finite-size scaling collapse of coercivity plateau. (b) Detailed scalings in the relatively fast-driving regimes after the plateau, where $W(v_{H})$ denotes the Lambert W function (real-valued branch). A bounded driving amplitude $H_{m}=1$ is imposed in (a), while the driving amplitude is unconstrained to ensure full saturation of the order parameter in (b). In this plot, we choose $\tau_{0}=2,\beta=2,J=1$.
Conclusion and discussion.—We have introduced the concept of the coercivity panorama to characterize the irreversibility of hysteresis phenomena in interacting systems across different time scales. In the periodically driven stochastic $\phi^{4}$ model, we analytically obtained the finite-time and finite-size scaling of the coercivity. The coercivity panorama bridges the gap between irreversibility in the linear response regime and irreversibility with signatures of FOPT, reflecting the competition between the thermodynamic limit and the quasi-static limit. Remarkably, the scaling relations near the coercivity plateau are universal, which is confirmed in the Curie-Weiss model. However, in the fast-driving regime, where the system transitions from the finite-time behavior of FOPT to the DPT regime, the stochastic $\phi^{4}$ and Curie-Weiss models exhibit different scaling behaviors.
We end this Letter with three remarks: i) Currently, the scaling exponents associated with hysteresis loops measured experimentally in different magnetic systems vary, and there are ambiguities in their theoretical explanations (See End Matter for details). The coercivity panorama we present provides a reliable basis for experimentally measuring the dynamic scaling of hysteresis loops. In particular, the finite-size scaling of coercivity is expected to be verified using microscopic experimental platforms such as cold-atom systems [66, 51] or two-dimensional ferromagnetic materials [82, 83]. ii) The phenomenological models used in the field of magnetism to study hysteresis loops and the research on the dynamics of FOPT in statistical physics previously seemed to be two independent areas. This work establishes a connection between them and presents a theoretical framework for coercivity based on microscopic dynamics. iii) The dynamics of FOPT and DPT can be uniformly characterized by coercivity. On this basis, the finite-time thermodynamics of interacting systems with phase transition driven at different time scales can be systematically studied. The optimal control of finite-time phase transitions [80, 54] with minimal energy cost, the performance of thermal machines with phase transition [84, 85, 86], and the information thermodynamics of interacting systems [87] are ̵fascinating issues worthy of further exploration.
Acknowledgments.— We thank Sai Li for the helpful discussions in the early stage of this work. This work is supported by the National Natural Science Foundation of China under grant No. 12305037 and the Fundamental Research Funds for the Central Universities under grant No. 2023NTST017.
End Matter
Appendix A: Dynamic hysteresis of Curie-Weiss model—The free energy of the Curie-Weiss model in an external field $H$ reads $\mathcal{F}_{H}(M)=-JM^{2}/(2N)-HM-\beta^{-1}\ln\Omega(M)$, where $M$ is the total magnetization taking values $-N,-N+2,\dots,N-2,N$. $N$ is the number of spins, $J$ is the interaction strength, $\Omega(M)$ is the number of microstates corresponding to a total magnetization $M$, $\beta\equiv 1/(k_{\mathrm{B}}T)$ is the inverse temperature, and $k_{\mathrm{B}}$ is the Boltzmann constant. The system exhibits spontaneous magnetization at zero field when $\beta>1/J$. The time evolution of the probability $P(M,t)$ for the system to have magnetization $M$ at time $t$ is governed by a master equation
$\frac{\partial P(M,t)}{\partial t}=\sum_{\eta=\pm}$ $\left[W_{\eta}(M_{-\eta},H)P(M_{-\eta},t)\right.$ (9)
$\left.-W_{\eta}(M,H)P(M,t)\right],$
with the transitions $M\to M_{\pm}\equiv M\pm 2$ occurring with rates,
$W_{\pm}(M,H)=\frac{N\mp M}{2\tau_{0}}\exp\left[\pm\beta\left(J\frac{M\pm 1}{N}+H\right)\right],$ (10)
where $\tau_{0}$ denotes the microscopic relaxation time associated with single spin flips. By numerically solving this master equation under a triangularly swept periodic external field $H(t)$, the coercivity can be extracted from the resulting hysteresis loops. Figures 4(a) and (b) show the hysteresis loops at various driving rates with $m\equiv M/N$. The extracted coercivities are shown in Fig. 4(c), demonstrating the four regimes investigated in the main text.
Figure 4: Hysteresis loops and coercivity panorama of the Curie-Weiss model with spin number $N=30$. Hysteresis loops at (a) $v_{H}=10^{-6}$(gray solid), $10^{-4}$(yellow dashed), $10^{-1}$(green dash-dotted); (b) $v_{H}=0.5$(solid), $1.0$(dashed), $2.0$(dash-dotted); (c) Coercivity $H_{c}$ as a function of driving rate $v_{H}$ with a limited driving amplitude $(H_{m}=1).$ The parameters are the same as Fig. 3.
To find out the correspondence between $\sigma$ in the stochastic $\phi^{4}$ model and $N$ in the Curie-Weiss model, we first derive the Fokker-Planck equation corresponding to Eq. (9) for $N\gg 1$ [69],
$\frac{\partial P(m,t)}{\partial t}=$ $-\frac{2+2\beta J/N}{\tau_{0}}\frac{\partial}{\partial m}\left[T(m,H)P(m,t)\right]$ (11)
$+\frac{1}{N}\frac{2}{\tau_{0}}\frac{\partial^{2}}{\partial m^{2}}\left[S(m,H)P(m,t)\right],$
where $T(m,H)=\sinh\beta(Jm+H)-m\cosh\beta(Jm+H)$ and $S(m,H)=\cosh\beta(Jm+H)-m\sinh\beta(Jm+H)$. Expanding Eq. (11) around the spinodal point $(m^{*},H^{*})$ and keeping the leading order of each term, the result reads
$\frac{\partial P(\hat{m},t)}{\partial t}\approx$ $-\frac{2}{\tau_{0}}\frac{\partial}{\partial\hat{m}}\left[\beta J\sqrt{\beta J-1}\hat{m}^{2}+\sqrt{\frac{\beta}{J}}\hat{H}\right]P(\hat{m},t)$ (12)
$+\frac{1}{N}\frac{2}{\tau_{0}}\frac{1}{\sqrt{\beta J}}\frac{\partial^{2}}{\partial\hat{m}^{2}}P(\hat{m},t),$
with $\hat{m}\equiv m-m^{*}$ and $\hat{H}\equiv H-H^{*}$. By comparing the diffusion coefficient in Eqs. (4) and that in the Fokker-Planck equation for the stochastic $\phi^{4}$ model, one obtains $\sigma^{2}\sim 1/N$.
Table 2: Experimental results of hysteresis scaling relations
System Scaling Methodology
Structural transition in VO2[75] $A\propto v_{H}^{1/2}$ $A\propto v_{H}^{\alpha}H_{m}^{\beta},\alpha=0.33\pm 0.03,\beta=0.66\pm 0.05,$ from a 3D continuous $(\Phi^{2})^{2}$ field theory with $O(N)$ symmetry in the $N\to\infty$ limit[52].
PbTiO3/polymer ferroelectric composites [29]
$A\propto v_{H}^{1/3}H_{m}^{2/3}$(saturated)
$A\propto v_{H}^{-1}H_{m}^{2}$(unsaturated)
Ferroelectric SrBi2Ta2O9 thin films [28]
$A\propto v_{H}^{2/3}H_{m}^{2/3}$(saturated)
$A\propto v_{H}^{-1/3}H_{m}^{2}$(unsaturated)
Ferroelectric BaTiO3 single crystals [27]
$A\propto v_{H}^{1.667H_{m}-2.804}H_{m}^{4.157}$(saturated)
$A\propto v_{H}^{-0.195}H_{m}^{0.95}$(unsaturated)
Ferroelectric BaTiO3 bulk ceramics [40]
$A\propto v_{H}^{-0.39}H_{m}^{1.06}$(saturated)
$A\propto v_{H}^{-0.55}H_{m}^{3.4}$(unsaturated)
Soft Pb(Ti, Zr)O3 ferroelectric ceramic [39]
$A\propto v_{H}^{-0.25}H_{m}^{0.89}$(saturated)
$A\propto v_{H}^{-1/3}H_{m}^{3}$(unsaturated)
Hard Pb(Ti, Zr)O3 ferroelectric ceramic [38]
$A\propto v_{H}^{-0.28}H_{m}^{0.89}$(saturated)
$A\propto v_{H}^{-0.43}H_{m}^{3.19}$(unsaturated)
Switching of bistable laser [23] $A-A_{0}\propto v_{H}^{2/3}$ $A-A_{0}\propto v_{H}^{n}$, $n$ approaches 2/3, predicted by the mean-field theory [23, 42]
Cold atomic system [66] $A-A_{0}\propto v_{H}^{0.64\pm 0.04}$
Mott transition in V2O3  [50]
$A-A_{0}\propto v_{H}^{0.62\pm 0.06}$(heating)
$A-A_{0}\propto v_{H}^{0.64\pm 0.09}$(cooling)
N-SmA transition in binary mixture [9]
$A-A_{0}\propto v_{H}^{0.672\pm 0.008}$(mixture ratio4:6:464:64 : 6)
$A-A_{0}\propto v_{H}^{0.766\pm 0.05}$(mixture ratio3:7:373:73 : 7)
$A-A_{0}\propto v_{H}^{0.701\pm 0.04}$(mixture ratio2:8:282:82 : 8)
$A-A_{0}\propto v_{H}^{0.629\pm 0.005}$(mixture ratio1:9:191:91 : 9)
Martensitic transition in Co [10]
$T_{p}-T_{0}\propto v_{H}^{0.39}$(heating)
$T_{p}-T_{0}\propto v_{H}^{0.49}$(cooling)
$T_{p}$ is the peak temperature,
$T_{0}$ is the characteristic reference temperature
Appendix B: Experimental results on rate-dependence of hysteresis loop area— Tab. 2 makes clear that the underlying protocols and observables vary widely. The $A-A_{0}\propto v_{H}^{2/3}$ scaling relation, as a universal prediction of mean‑field theories [42], is observed experimentally in systems such as the Mott transition in $V_{2}O_{3}$ [50], switching of bistable laser [23]. Furthermore, exponents in $A\propto v_{H}^{\alpha}$, close to 1/3 [29] and 1/2 [75], are explained by our current work, further validating the applicability of our theory. Besides these typical scalings, a broader range of exponent behavior has been observed across systems, especially since many scalings break down outside narrow frequency. For example, in certain ferroelectric ceramics [40, 39, 38], the measurements fall in the high-frequency regime and exhibit negative exponents. The difference of the exponents stems from variations in amplitude $H_{m}$, observables ($A$ or $A-A_{0}$), and whether measurements probe saturated or unsaturated regimes by tuning $v_{H}$ or $H_{m}$. Moreover, the poorly defined “quasi-static” plateau height ($H_{p}$ or $A_{0}$) systematically skews the fitted exponent. In practice, the true plateau is often unreachable and replaced by the minimal observed coercivity or an arbitrary fit constant. However, since their actual experimental conditions might not strictly meet the slow-drive FOPT regime for 2/3 scaling, the determined $A_{0}$ may not accurately reflect the true quasi-static limit. This discrepancy occurs because, without a panoramic view, researchers can only compare their results with existing slow-driving theories. By mapping all data onto a single, unified control variable, our coercivity landscape overcomes system-dependent biases and reconciles these disparate results. It also offers a coherent framework to explain the anomalous scaling trends that resisted interpretation under previous microscopic or empirical models.
Supplemental Materials for:
“Coercivity Panorama of Dynamic Hysteresis”
This appendix is devoted to providing the detailed derivations and the supporting discussions to the main text. The contents are listed as follows.
Appendix SI The Stochastic $\phi^{4}$ Model
In this section, we present detailed explanations and derivations for scaling relations in the stochastic $\phi^{4}$ model.
SI.1 Potential landscape, Dynamic equation and Hysteresis curves
To investigate the scaling behavior of coercivity in detail, we adopt the $\phi^{4}$ model to describe a field-driven first-order phase transition (FOPT). To capture the finite-size effects in specific models, a noise term is introduced into the system’s dynamics to represent finite fluctuations and the resulting non-deterministic behavior away from the thermodynamic limit [68, 52].
Substituting the free energy expression [Eq. (1) in the main text] into the dynamic equation [Eq. (2) in the main text], the specific form of the dynamic equation reads,
$\frac{\partial\phi}{\partial t}=-\lambda a_{2}\phi-\lambda a_{4}\phi^{3}+\lambda H+\zeta(t).$ (13)
The change of the order parameter over a small time interval $\delta t$ is given by the discretized form of the above equation,
$\delta\phi=-\lambda a_{2}\phi\delta t-\lambda a_{4}\phi^{3}\delta t+\lambda H\delta t+\sqrt{2\lambda\sigma^{2}\delta t}W,$ (14)
where $W$ is a random variable drawn from a standard normal distribution. By simulating Eq. (14), we obtain the evolution of $\phi(t)$ under varying field $H(t)$.
Figure 5: Potential curve of the $\phi^{4}$ model at different external field value.
The three subfigures in Fig. 5 show the potential curves at different values of the external field. From left to right, the field is strongly negative, zero, and strongly positive, respectively. We control the external field increasing linearly from a strongly negative to a strongly positive value, with the system initially prepared in the equilibrium state corresponding to the strongly negative field. As the field changes, the averaged order parameter $\langle\phi\rangle$ evolves from a negative to a positive value, and the point at which it crosses zero defines the coercivity. We record the coercivity for different driving velocities $v_{H}$ and noise strengths $\sigma$.
Figure 6: Hysteresis curve for $\sigma=0.1$ (solid lines) and $\sigma=1.0$ (dased lines). (a) Slow driving. (b) Fast driving.
Figures 6(a) and 6(b) show the field-order parameter ($H-\langle\phi\rangle$) curves during relatively slow and fast driving processes, respectively. The coercivity, as reflected in the $\langle\phi\rangle-H$ curves, increases with the driving velocity. In the relatively slow driving regime, the curves exhibit distinct behaviors depending on the noise strength: for large noise strength, the transitions are smoother and the coercivity is smaller; for small noise strength, the transitions are sharper and the coercivity is larger. This difference arises from the noise-induced reduction in relaxation time. As the external field increases and the left potential well becomes metastable, the escape rate to the stable well is enhanced by larger noise strength [88]. As a result, the order parameter transitions earlier, at a smaller field value. In the relatively fast driving regime, the $H-\langle\phi\rangle$ curves for different noise strengths nearly overlap. This is because the driving velocity exceeds any intrinsic relaxation timescale of the system, regardless of the noise strength, leading to similar passive relaxation processes.
SI.2 $H_{c}\sim v_{H}$ scaling near equilibrium
Linear scaling is observed in the near-equilibrium regime. In the equilibrium limit, the order parameter $\langle\phi\rangle$ becomes zero at $H=0$, indicating that the coercivity vanishes in this limit. Since the potential is symmetric with a double-well structure at $H=0$, the skewness of the equilibrium distribution, represented by the third central moment $\mu_{3}$, also vanishes at this point. In other words, both the equilibrium-limit values of $H_{c}$ and $\mu_{3}$ at $\langle\phi\rangle=0$ are zero.
To quantify deviations from equilibrium in the near-equilibrium regime, the key quantities, $H_{c}$, $\evaluated{\mu_{3}}_{\langle\phi\rangle=0}$, and $\evaluated{\left(d\langle\phi\rangle/dH\right)}_{\langle\phi\rangle=0}$, can be expanded in power series around zero,
$H_{c}=h_{1}v_{H}^{\alpha}+h_{2}v_{H}^{2\alpha}+\dots$ (15a)
$\mu_{3}=u_{1}v_{H}^{\beta}+u_{2}v_{H}^{2\beta}+\dots$ (15b)
$\evaluated{\frac{d\langle\phi\rangle}{dH}}_{\langle\phi\rangle=0}=\evaluated{\frac{d\langle\phi\rangle_{\mathrm{eq}}}{dH}}_{\langle\phi\rangle_{\mathrm{eq}}=0}+g_{1}v_{H}^{\gamma}+g_{2}v_{H}^{2\gamma}+\dots,$ (15c)
where the subscript $\mathrm{eq}$ denotes equilibrium limit. Substituting the above series into Eq. (6) in the main text and retaining of the lowest-order corrections in each term, we find that $\alpha=\beta=1$ in order to balance the powers of $v_{H}$. Therefore, the near-equilibrium correction of the coercivity follows as
$H_{c}\propto v_{H}.$ (16)
We emphasize that the above analysis applies to finite noise strength, as $\evaluated{\left(d\langle\phi\rangle/dH\right)}_{\langle\phi\rangle=0}$ diverges in the equilibrium limit when the noise vanishes, and thus cannot be expanded in a power series like Eq. (15c) near equilibrium.
SI.3 $Hc\sim v_{H}^{1/2}$ scaling in the fast-driving limit
In the fast-driving regime, coercivity shows no significant dependence on noise strength or system size. This is because the external field changes so rapidly that the system remains effectively frozen before the potential landscape becomes monostable. In this regime, the dynamics are dominated by deterministic forces, while fluctuations, which are crucial for transitions between metastable states, play a negligible role. Consequently, the coercivity behavior can be understood by analyzing the relaxation dynamics toward the monostable state in systems with vanishing fluctuations (or equivalently, in the thermodynamic limit) under a time-dependent strong field.
For a varying external field, the position of the monostable state, that is the equilibrium state, in the $\phi^{4}$ model also changes accordingly. When the magnitude of the field is sufficiently large, specifically for $|H|\gg\sqrt{|a_{2}^{3}/a_{4}|}$, the equilibrium state can be approximated as follows
$\phi_{\mathrm{eq}}=\left(\frac{H}{a_{4}}\right)^{1/3},$ (17)
In deriving the above expression, the $a_{2}\phi$ term in $\partial f(\phi,H)/\partial\phi$ has been neglected, as it becomes negligible compared to the $a_{4}\phi^{3}$ term when the external field $H$ is sufficiently large.
The relaxation dynamics near the equilibrium state can be approximated by the following linearized equation,
$\frac{\partial\langle\phi\rangle}{\partial t}\approx-\lambda\evaluated{\frac{\partial^{2}f(\phi,H)}{\partial\phi^{2}}}_{\phi_{\mathrm{eq}}}\left(\langle\phi\rangle-\phi_{\mathrm{eq}}\right),$ (18)
from which we obtain that the characteristic relaxation time to the equilibrium state is
$\tau_{\mathrm{relax}}=$ $\frac{1}{\lambda\evaluated{(\partial^{2}f/\partial\phi^{2})}_{\mathrm{eq}}}=\frac{1}{\lambda(a_{2}+3a_{4}\phi_{\mathrm{eq}}^{2})}$ (19)
$\approx$ $\frac{1}{3\lambda a_{4}\phi_{\mathrm{eq}}^{2}}$
$\approx$ $\frac{1}{3\lambda a_{4}^{1/3}H^{2/3}},$
where we have used the expression of $\phi_{\mathrm{eq}}$.
With an increasing external field, the equilibrium state moves with the rate
$r_{\phi_{\mathrm{eq}}}\equiv\frac{d\phi_{\mathrm{eq}}}{dH}v_{H}\approx\frac{1}{3}\frac{v_{H}}{a_{4}^{1/3}H^{2/3}}.$ (20)
The characteristic rate of relaxation is
$r_{\mathrm{relax}}\equiv\frac{1}{\tau_{\mathrm{relax}}}\approx 3\lambda a_{4}^{1/3}H^{2/3}.$ (21)
Since the characteristic rate of relaxation increases and the moving rate of the equilibrium state decreases with the external field for fixed $v_{H}$, we identify that the equation
$r_{\mathrm{relax}}\simeq r_{\phi_{\mathrm{eq}}}$ (22)
represents a turning point of the field that the relaxation of the system becomes able to catch up the variation of the equilibrium state and we define this turning point as $H_{\mathrm{satur}}$. By substituting the expressions of $r_{\mathrm{relax}}$ and $r_{\phi_{\mathrm{eq}}}$ into the above equation, we obtain $H_{\mathrm{satur}}$ for the $\phi^{4}$ model scales with $v_{H}$ as follows,
$H_{\mathrm{satur}}\propto v_{H}^{3/4}.$ (23)
Figure 7: Geometric method to explain the scaling exponent in the fast driving regime.
To characterize the varying scaling exponent of coercivity with respect to the driving velocity in the fast-driving regime, we introduce a reference right triangle on the hysteresis curve. The hypotenuse of the triangle is defined by the line passing through the coercivity point with the same slope as the $H$-$\langle\phi\rangle$ curve at this point, and it intersects the curve again at its terminal point. The vertical leg of the triangle is defined to pass through the initial point of the hysteresis curve, and the horizontal leg is determined accordingly.
Since the field value of the left terminal point of the horizontal leg is much smaller than that of the right terminal point, and the averaged order parameter value approaches the equilibrium value at the right terminal point, the length of the horizontal leg is approximated by the magnitude of $H_{\mathrm{satur}}$. Acoordingly, the length of the vertical leg is approximated by $\phi_{\mathrm{eq}}(H_{\mathrm{satur}})$. Then the slope of the hypotenuse is approximated by $\phi_{\mathrm{eq}}(H_{\mathrm{satur}})/H_{\mathrm{satur}}$, which gives the following scaling relation with respect to $v_{H}$,
$\evaluated{\frac{d\langle\phi\rangle}{dH}}_{\langle\phi\rangle=0}\propto H_{\mathrm{satur}}^{-2/3}\propto v_{H}^{-1/2}.$ (24)
Furthermore, in the fast driving regime, the skewness of $\phi$ is neglible, and the coercivity is approximated by
$H_{c}=v_{H}\evaluated{\frac{d\langle\phi\rangle}{dH}}_{\langle\phi\rangle=0},$ (25)
which is verified in Fig. 6(b), where the tilted dashed line is plotted with the slope $H_{c}/v_{H}$. Hence, the scaling exponent of the coercivity with respect to dirving velocity is $1/2$ in the fast driving regime, that is
$H_{c}\propto v_{H}^{1/2}.$ (26)
SI.4 Finite-size scaling of the FOPT plateau
In this section, we investigate how finite-size effects influence the coercivity plateau characteristic of first-order transitions. By expanding the potential near the spinodal point, we derive how the plateau height and width scale with system size.
Since the plateau is associated with the first-order phase transition, we will expand the quartic potential near the transition point defined in the thermodynamic limit, retaining terms up to the cubic order of the order parameter [23, 53]. This is justified by the fact that the order parameter undergoes a discontinuous flip during the first-order transition. The transition point of the system in the thermodynamic limit is exactly the spinodal point $(H_{\mathrm{sp}},\phi_{\mathrm{sp}})$ of the potential $f(\phi,H)$, which is determined by $\partial f/\partial\phi=0$ and $\partial^{2}f/\partial\phi^{2}=0$,
$H_{\mathrm{sp}}=\pm\frac{2}{3}a_{2}\sqrt{\frac{-a_{2}}{3a_{4}}},\quad\phi_{\mathrm{sp}}=\pm\sqrt{\frac{-a_{2}}{3a_{4}}}.$ (27)
Defining $h\equiv H-H_{\mathrm{sp}}$ and $\varphi\equiv\phi-\phi_{\mathrm{sp}}$, the original quartic potential is expanded as
$f(\varphi,h)=\frac{1}{3}a_{3}\varphi^{3}-h\varphi,$ (28)
where $a_{3}=\pm\sqrt{-3a_{2}a_{4}}$. Substitute the above cubic potential into the governing Langevin equation and then one obtains the corresponding discretized equation as follows,
$\delta\varphi=-a_{3}\varphi^{2}v_{H}^{-1}\delta h+hv_{H}^{-1}\delta h+\sqrt{2\sigma^{2}v_{H}^{-1}\delta h}W,$ (29)
where we have used the relation $\delta h=\lambda v_{H}\delta t$ for linearly varying field $H(t)=H_{0}+\lambda v_{H}t$. In Ref. [53], Zhong applies scaling transformations to the dynamic equation governed by the cubic potential, based on renormalization-group theory, and derives complete scaling relations for all system parameters and variables with respect to the driving velocity. In the present work, to investigate the finite-size effect on the coercivity plateau, we perform analogous scaling transformations with respect to the noise strength. Using $\Sigma$ as the scaling factor to scale the noise strength from $\sigma$ to $\sigma^{\prime}=\sigma\Sigma^{-1}$, the quantity $\mathcal{O}$ becomes $\mathcal{O}^{\prime}=\mathcal{O}\Sigma^{-[\mathcal{O}]}$, where $[\mathcal{O}]$ is the scaling dimension of $\mathcal{O}$. Performing scaling transformations on Eq. (29), one obtains
$\delta\varphi^{\prime}\Sigma^{[\varphi]}=$ $-a_{3}^{\prime}\varphi^{\prime 2}v_{H}^{\prime-1}\delta h^{\prime}\Sigma^{[a_{3}]+2[\varphi]-[v_{H}]+[h]}$ (30)
$+h^{\prime}v_{H}^{\prime-1}\delta h^{\prime}\Sigma^{2[h]-[v_{H}]}$
$+\sqrt{2\sigma^{\prime 2}v_{H}^{\prime-1}\delta h^{\prime}}\Sigma^{1-[v_{H}]/2+[h]/2}.$
To keep the dynamic equation invariant after scaling transformation, the scaling dimensions of each term in the equation should balance, that is
$[\varphi]=[a_{3}]+2[\varphi]-[v_{H}]+[h]=2[h]-[v_{H}]=1-[v_{H}]/2+[h]/2.$ (31)
According to renormalization-group theory, the scaling dimension of the coefficient in the highest-order term is set to be zero[53], that is $[a_{3}]=0$. Substituting $[a_{3}]=0$ into Eq. (31), we obtain that
$[h]=\frac{4}{3},\quad[\varphi]=\frac{2}{3},\quad[v_{H}]=2.$ (32)
The above scaling dimensions reveal that the scaling exponents of $h$, $\varphi$ and $v_{H}$ with respect to $\sigma$ are respectively $4/3$, $2/3$ and $2$. These scaling exponents are applicable for cases of small noise strengths and slow driving velocities, where the jump of first-order phase transition is prominent. Since the coercivity plateau is also a feature of first-order phase transition, we conclude the scaling exponents as
$|H_{c}^{*}-H_{\mathrm{sp}}|\propto\sigma^{4/3}\quad\textbf{at}\quad v_{H}^{*}\propto\sigma^{2},$ (33)
where $H_{c}^{*}$ is the height of the coercivity plateau and $|H_{c}^{*}-H_{\mathrm{sp}}|$ is the difference in the height between systems with finite and vanishing noise strengths. Although these scaling relations are directly applicable to turning points like the spinodal point in the hysteresis loops, it is verified that the coercivity points also possess these scaling relations due to the small difference in turning point and coercivity point for rapid transition process. Considering that the plateau height decreases with increasing $\sigma$, as shown in Fig. 8(a), we rewrite the above relations as
$H^{*}=H_{\mathrm{sp}}-A\sigma^{4/3}\quad\textbf{at}\quad v_{H}^{*}\propto\sigma^{2},$ (34)
where $A$ is a constant and $v_{H}^{*}$ is the reference velocity to observe coercivity plateau.
SI.5 Reference point of the FOPT plateau
Fig. 8(a) is redrawn from Fig. 1(f) in the main text. We observe that the FOPT plateau becomes broader and higher with decreasing noise (increasing size) strength. The specific FOPT plateau point is defined as the local minimum point shown in Fig. 8(b), while the plateau point for very small noise strength has not been achieved in the presented simulation timescale.
Figure 8: Coercivity and scaling exponent at different driving velocity for different noise strengths.
SI.6 $H_{c}-H_{P}\sim(v_{H}-v_{P})^{2/3}$ scaling in the post-plateau regime
Acoording to Eq. (5) in the main text, for a given initial condition, the hysteresis curve under linearly varying field is uniquely determined by the driving rate $v_{H}$. Considering the specific hysteresis curve corresponding to $v_{P}$ for some $\sigma$, denote the turning point within this curver as $(H_{P}^{*},\phi_{P}^{*})$. Expanding the free energy around this turning point and retaining to the third order term,
$f_{3}(\varphi_{P},h_{P})=b_{0}+b_{1}\varphi_{P}+\frac{1}{2}b_{2}\varphi_{P}^{2}+\frac{1}{3}b_{3}\varphi_{P}^{3}-\phi_{P}^{*}h_{P}-h_{P}\varphi_{P},$ (35)
with $\varphi_{P}\equiv\phi-\phi_{P}^{*}$, $h_{P}\equiv H-H_{P}^{*}$, and
$b_{0}=\frac{1}{4}a_{4}\phi_{P}^{*4}+\frac{1}{2}a_{2}\phi_{P}^{*2}-H_{P}^{*}\phi_{P}^{*},$
$b_{1}=a_{4}\phi_{P}^{*3}+a_{2}\phi_{P}^{*}-H_{P}^{*},$
$b_{2}=3a_{4}\phi_{P}^{*2}+a_{2},$
$b_{3}=3a_{4}\phi_{P}^{*}.$
Substituting Eq. (35) into the governing Fokker-Planck equation,
$\frac{\partial P(\varphi_{P},t)}{\partial t}=\lambda\frac{\partial}{\partial\varphi_{P}}\left[\left(b_{1}+b_{2}\varphi_{P}+b_{3}\varphi_{P}^{2}-h_{P}\right)P(\varphi_{P},t)\right]+\lambda\sigma^{2}\frac{\partial^{2}P(\varphi_{P},t)}{\partial\varphi_{P}^{2}}.$ (36)
Multiplying both sides of the above equation by $\varphi_{P}$ and integrating over $\varphi_{P}$, one obtains the averaged equation
$v_{H}\frac{d\langle\varphi_{P}\rangle}{dh_{P}}=-b_{1}-b_{2}\langle\varphi_{P}\rangle-b_{3}\langle\varphi_{P}^{2}\rangle+h_{P},$ (37)
where the relation $dh_{P}=\lambda v_{H}dt$ has been used. Evaluating the above equation at $\langle\varphi_{P}\rangle=0$, the result reads
$v_{H}\evaluated{\frac{d\langle\varphi_{P}\rangle}{dh_{P}}}_{\langle\varphi_{P}\rangle=0}=-b_{1}-b_{3}\evaluated{\langle\varphi_{P}^{2}\rangle}_{\langle\varphi_{P}\rangle=0}+\evaluated{h_{P}}_{\langle\varphi_{P}\rangle=0}.$ (38)
For $v_{H}=v_{P}$, the above relation becomes
$v_{P}\evaluated{\frac{d\langle\varphi_{P}\rangle}{dh_{P}}}_{\langle\varphi_{P}\rangle=0,v_{P}}=-b_{1}-b_{3}\evaluated{\langle\varphi_{P}^{2}\rangle}_{\langle\varphi_{P}\rangle=0,v_{P}}.$ (39)
Defining $\nu_{h}\equiv v_{H}-v_{P}$ and expanding Eq. (38) around $\nu_{h}=0$, we obtain
$\nu_{h}s=-b_{3}\mu+\evaluated{h_{P}}_{\langle\varphi_{P}\rangle=0},$ (40)
where $s\equiv\evaluated{(d\langle\varphi_{P}\rangle/dh_{P})}_{\langle\varphi_{P}\rangle=0}-\evaluated{(d\langle\varphi_{P}\rangle/dh_{P})}_{\langle\varphi_{P}\rangle=0,v_{P}}$, and $\mu\equiv\evaluated{\langle\varphi_{P}^{2}\rangle}_{\langle\varphi_{P}\rangle=0}-\evaluated{\langle\varphi_{P}^{2}\rangle}_{\langle\varphi_{P}\rangle=0,v_{P}}$. To determine the leading order on the left side of the above equation, we perform scaling transformations with respect to $\nu_{P}$, which gives $[\nu_{P}]=1$, $[s]=[\varphi_{P}]-[h_{P}]$, $[\mu]=2[\varphi_{P}]$, and set $[b_{3}]=0$. The resulting scaling relation is $\evaluated{h_{P}}_{\langle\varphi_{P}\rangle=0}\sim\nu_{P}^{2/3}$. Due to the closeness between the turning point and the coercivity point for hysteresis loops with a feature of first-order phase transition, the coercivity value also adheres to this scaling, that is $H_{c}-H_{P}\sim(v_{H}-v_{P})^{2/3}$.
Appendix SII The Curie-Weiss model
In this section, we provide detailed dynamic equations and derivations of the scaling relations for the Curie-Weiss model.
SII.1 Dynamic equations
To validate and extend the scaling behavior of coercivity in a different microscopic context, we analyze the Curie-Weiss model under dynamic field driving. We reveal two distinct scaling regimes in the Curie-Weiss model: $H_{c}-H^{*}\propto v_{H}^{2/3}$, arising from excess work and delayed transitions from metastable states, and $H_{c}\propto v_{H}^{1/3}$, derived from Airy-function analysis near the critical point. In the strong-field limit, a transcendental relation involving the Lambert W-function further characterizes the scaling of coercivity.
The Curie-Weiss (C-W) model is a paradigmatic mean-field model for ferromagnetic systems, capturing phase transitions through infinite-range interactions. Consider a system of $N$ Ising spins $\sigma_{i}=\pm 1$, labeled $i=1,...,N$, where each spin interacts as the total magnetization with all others via a ferromagnetic coupling $J/2N$, and define $M=\sum_{i=1}^{N}\sigma_{i}$. The system is immersed in a thermal bath at inverse temperature $\beta=1/(k_{B}T)$ and driven by an external time-dependent magnetic field $H(t)$ The Hamiltonian is given by [79, 54]:
$\mathcal{H}=-\frac{J}{2N}\left(\sum_{i=1}^{N}\sigma_{i}\right)^{2}-H(t)\sum_{i=1}^{N}\sigma_{i},$ (41)
When an arbitrary spin flips from $\mp 1$ to ±$\pm 1$, the total magnetization will change by two units, i.e., $M\to M_{\pm}\equiv M\pm 2$. Following the master equation formulation and transition rates from stochastic dynamics, the probability $P(M,t)$ of observing total magnetization $M$ evolves as:
$\frac{dP(M,t)}{dt}=\sum_{\eta=\pm}\left[W_{\eta}(M-\eta,H)P(M-\eta,t)-W_{\eta}(M,H)P(M,t)\right],$ (42)
with the transition rates for $M\to M_{\pm}\equiv M\pm 2$,
$W_{\pm}(M,H)=\frac{N\mp M}{2\tau_{0}}\exp\left[\pm\beta\left(J\frac{M\pm 1}{N}+H\right)\right].$ (43)
These transition rates satisfy the detailed balance condition, ensuring microscopic reversibility in equilibrium:
$W_{\pm}(M,H)P_{\text{eq}}(M,H)=W_{\mp}(M_{\pm},H)P_{\text{eq}}(M_{\pm},H),$ (44)
where $P_{\text{eq}}(M,H)\propto\exp\left[-\beta\mathcal{F}_{H}(M)\right]$ is the equilibrium distribution with the free energy
$\mathcal{F}_{H}(M)=-\frac{J}{2N}M^{2}-HM-\frac{1}{\beta}\ln\Omega(M),$ (45)
and $\Omega(M)=\frac{N!}{[(N+M)/2]![(N-M)/2]!}$ is the microscopic degeneracy of magnetization $M$. The master equation can be expressed in the matrix form
[…]
where $W_{-}(-N,H)=W_{+}(N,H)=0$ by definition.
To obtain the time evolution of the average magnetization $\langle M\rangle$, we multiply both sides of the master equation by $M$ and sum over all possible states.
$\frac{d\langle M\rangle}{dt}=\sum_{M}\sum_{\eta=\pm}M\left[W_{\eta}(M-\eta,H)P(M-\eta,t)-W_{\eta}(M,H)P(M,t)\right].$ (47)
Substituting the explicit form of the transition rates, we arrive at a more detailed expression for the time derivative of $\langle M\rangle$:
$\frac{d\langle M\rangle}{dt}=\sum_{M}\Bigg{[}$ $\frac{N-M+2}{2\tau_{0}}e^{\beta J(M-1)/N+\beta H}P(M-2,t)+\frac{N+M+2}{2\tau_{0}}e^{-\beta J(M+1)/N-\beta H}P(M+2,t)$
$-\left(\frac{N-M}{2\tau_{0}}e^{\beta J(M+1)/N+\beta H}+\frac{N+M}{2\tau_{0}}e^{-\beta J(M-1)/N-\beta H}\right)P(M,t)\Bigg{]}M.$ (48)
with the microscopic relaxation time $\tau_{0}$.
In the thermodynamic limit $N\to\infty$, the fluctuations become negligible and the system’s state is characterized by the mean magnetization $m=M/N$. The Hamiltonian of the system can be expressed as
$\mathcal{H}(m)=-\frac{J}{2}m^{2}-H(t)m.$ (49)
Defining the per‐spin rates
$w_{\pm}(m)=\lim_{N\to\infty}\frac{1}{N}W_{\pm}(Nm)=\frac{1\mp m}{2\,\tau_{0}}\,\exp\!\bigl{[}\pm\beta\,(Jm+H)\bigr{]},$ (50)
we arrive at the deterministic evolution equation
$\frac{dm}{dt}=\frac{2}{\tau_{0}}[\sinh{(\beta Jm+\beta H)}-m\cosh{(\beta Jm+\beta H)}]=g(m,H)$ ) (51)
We analyze a scenario where the external magnetic field $H(t)$ evolves from an initial value $H_{i}$ to a final value $H_{f}$ via a linear schedule:
$H(t)=H_{i}+v_{H}t.$ (52)
Here, the quench rate is defined by $v_{H}=(H_{f}-H_{i})/t_{f}$, with $t_{f}$ representing the total duration of the process. To visualize dynamic hysteresis loops in the Curie-Weiss model, we numerically solve the master equation Eq.(42) and track the average magnetization $\langle m\rangle$ as a function of $H(t)$. For finite systems, the total magnetization $M$ is treated stochastically, with its probability distribution $P(M,t)$ evolving according to the master equation. The per-site average $\langle m\rangle=\sum_{M}MP(M,t)/N$ is computed from this distribution.
SII.2 $2/3$ power law scaling of coercivity
During the dynamic process described by Eq. (51), the system evolves along trajectories that may adhere to local equilibrium in either metastable or globally stable states, until a sudden transition occurs into a distinct state. This transition, contingent on the system’s history, exemplifies a finite-time first-order phase transition, driven by the collapse of the metastable state. The turning points (±$(\pm H^{*},\pm m^{*})$ are solved from $g(m,H)=0$ and $\partial_{m}g(m,H)=0$ as
$m^{*}=-\sqrt{1-\frac{1}{\beta J}},H^{*}=\frac{1}{\beta}\left[\sqrt{\beta J(\beta J-1)}-\operatorname{arctanh}\left(\sqrt{\frac{\beta J-1}{\beta J}}\right)\right]$ (53)
Notably, when the quench occurs at a finite rate, the system’s dynamics lag behind the quasistatic evolution, giving rise to an energy discrepancy quantified by the excess work. Defined as $w_{ex}=w-w_{qs}$, the excess work between dynamic and quasistatic hysteresis loops quantifies the additional energy cost from non-equilibrium dynamics, highlighting irreversibility during phase transitions. The $v_{H}^{2/3}$ scaling relation for the excess work,
$w_{\text{ex}}\approx-\frac{A_{1}(1+\sqrt{1-\frac{1}{\beta J}})}{[4\sqrt{\beta J(\beta J-1)}\frac{\beta}{\tau_{0}^{2}}]^{1/3}}v^{2/3},$ (54)
is demonstrated in [54], which extends to various systems undergoing first-order phase transitions, including open quantum systems [89, 90, 91] and chemical reaction networks [92, 93, 94]. The results above are due to neglecting the fluctuations that vanish in infinitely large systems.When the quench rate is slow but not slower than the system’s relaxation rate in large systems, nonergodicity arises, and the quasistatic work can be approximated by the deterministic limit work. It is also observed that the average excess work scales as $w_{\text{ex}}\sim v^{2/3}$ when the quench rate is slow but not excessively slow in relatively large systems[54].
Notably, for systems not extremely small, in a slow quench, the excess work can be approximated by the enclosed rectangular area bounded by the upper and lower bounds of magnetization $m$ and the interval between $H_{c}$ and $H^{*}$,
$w_{\text{ex}}\approx\Delta m(H_{c}-H^{*}),$ (55)
where the magnetization difference across the transition $\Delta m\approx 2$ isn’t related to $v_{H}$. This leads to the scaling relation
$H_{c}-H^{*}\sim v_{H}^{2/3},$ (56)
which is numerically verified in the Fig. 9. It can be found that the finite-size system can only locally observe 2/3 scaling, and the smaller the system, the greater the deviation.
Figure 9: Coercivity delay behavior of the Curie-Weiss model as a function of driving rate with parameters $\tau_{0}=2,\beta=2,J=1$. Purple dots are obtained by numerically solving the master equation Eq.(42) in $N=30,150$. Gray dots are obtained by numerically solving the deterministic equation Eq.(51) in $N\to\infty$.
SII.3 $H_{c}\sim v^{1/3}_{H}$ scaling between FOPT plateau and fast-driving limit
While the potential landscape transitions from bistable to monostable at $t^{*}$, finite-rate quenches induce dynamical lag: the system delays switching to the global minimum by $\hat{t}_{\text{del}}$,defined via $m(t^{*}+\hat{t}_{\text{del}})=m^{*}$. Defining the shifted variables $\hat{m}=m-m^{*}$, $\hat{H}=H-H^{*}$, and $\hat{t}=t-t^{*}$, we derive by expanding the equation around the point $(m^{*},H^{*})$:
$\frac{d\hat{m}}{d\hat{t}}=A\hat{m}^{2}+B\hat{t},$ (57)
where
$A=\frac{2\beta J\sqrt{\beta J-1}}{\tau_{0}},\quad B=\frac{2v}{\tau_{0}}\sqrt{\frac{\beta}{J}}.$ (58)
To isolate the universal structure of the above equation, we introduce dimensionless variables via
$u=\alpha\,\hat{m},\qquad s=\gamma\,\hat{t},$ (59)
with the two scale factors $\alpha,\gamma>0$ chosen to eliminate all prefactors in the resulting equation. It can be adimensionalized into
$\frac{du}{ds}=\frac{A}{\alpha\,\gamma}\;u^{2}+\frac{\alpha\,B}{\gamma^{2}}\;s.$ (60)
To rescale the coefficients to unity, we have
$\frac{A}{\alpha\,\gamma}\;=1,\quad\frac{\alpha\,B}{\gamma^{2}}\;=1,$ (61)
which leads to the following expressions,
$\alpha$ $=A^{2/3}B^{-1/3}=\left(\frac{2}{\tau_{0}}\beta J\sqrt{\beta J-1}\right)^{2/3}\left(\frac{2}{\tau_{0}}\sqrt{\frac{\beta}{J}}v\right)^{-1/3},$ (62)
$\gamma$ $=A^{1/3}B^{1/3}=\left(\frac{2}{\tau_{0}}\beta J\sqrt{\beta J-1}\right)^{1/3}\left(\frac{2}{\tau_{0}}\sqrt{\frac{\beta}{J}}v\right)^{1/3}.$
With these rescaled coefficients, the original equation simplifies to the standard Riccati form
$\frac{du}{ds}=u^{2}+s,$ (63)
Under the asymptotic boundary condition as $s\to-\infty$, $u\to\sqrt{-s}$, the solution to the above equation is the Airy function
$u(s)=\frac{\mathrm{Ai}^{\prime}(-s)}{\mathrm{Ai}(-s)},$ (64)
Thus, the delay time is given by
$\hat{t}_{\text{del}}=-\mathrm{A^{\prime}_{1}}\left[4\sqrt{\beta J(\beta J-1)}\frac{\beta}{\tau_{0}^{2}}v_{H}\right]^{-1/3},$ (65)
with the first zero point of the Airy prime function $\mathrm{A^{\prime}_{1}}$. As shown in Fig.10(a) and (b), numerical observation suggests that the slope at the point $(m^{*},H^{*}+v_{H}\hat{t}_{\text{del}})$ corresponding to the delay time $\hat{t}_{\text{del}}$ is approximately equal to the slope connecting the coercivity and (0, -1), namely,
$\frac{dm}{dH}|_{H=H*}=\frac{\frac{d\hat{m}}{dt}}{v}|_{t=\hat{t}_{\mathrm{del}},\hat{m}=0}=\frac{1}{H_{c}}.$ (66)
We further conduct a detailed numerical verification of this result, as illustrated in Fig.10(c). By matching to the Airy‐function solution Eq.(65), the local slope at $\hat{m}=0$ is obtained as
$\frac{d\hat{m}}{d\hat{t}}|_{\hat{m}=0,\hat{t}=\hat{t}_{\text{del}}}=-\frac{2\mathrm{A^{\prime}_{1}}}{\tau_{0}}\sqrt{\frac{\beta}{J}}\left[4\sqrt{\beta J(\beta J-1)}\frac{\beta}{\tau_{0}^{2}}\right]^{-1/3}v_{H}^{2/3}$ (67)
Figure 10: Subplot (a) and (b) show numerical observation of the slope at $(m^{*},H^{*}+v_{H}\hat{t}_{\text{del}})$ and the slope connecting coercivity with (0, -1). Subplot (c) details the verification process. Here, we also choose $N=30$, $\tau_{0}=2,\beta=2,J=1$,$H_{m}=1$.
Rearranging the above equations results in the scaling relation
$H_{c}=-\frac{\tau_{0}}{2\mathrm{A^{\prime}_{1}}\sqrt{\beta J}}\left[4\sqrt{\beta J(\beta J-1)}\frac{\beta}{\tau_{0}^{2}}\right]^{1/3}v_{H}^{1/3}$ (68)
SII.4 $H_{c}$ scaling as the Lambert W function in the fast-driving limit
In the strong-field limit where $|H|\gg|m|$, the interaction term $Jm$ becomes negligible compared to $H$, and the deterministic mean-field equation simplifies to
$\frac{dm}{dt}\approx\frac{2}{\tau_{0}}\left[\sinh(\beta H)-m\cosh(\beta H)\right].$ (69)
This linear differential equation describes exponential relaxation toward the equilibrium value
$m_{\mathrm{eq}}\approx\tanh(\beta H)\approx 1-2e^{-2\beta H},\quad\text{for }\beta H\gg 1.$ (70)
Inverting the relation provides a useful approximation for the external field in terms of the equilibrium magnetization:
$H\approx-\frac{1}{2\beta}\ln(1-m_{\mathrm{eq}}).$ (71)
To estimate the characteristic relaxation time toward equilibrium, we linearize the dynamical equation around $m_{\mathrm{eq}}$:
$\frac{dm}{dt}\approx-\frac{2}{\tau_{0}}\cosh(\beta H)\,(m-m_{\mathrm{eq}}).$ (72)
This leads to the relaxation time
$\tau_{\mathrm{relax}}\approx\left(\frac{2}{\tau_{0}}\cosh(\beta H)\right)^{-1}\approx\frac{\tau_{0}}{2}e^{-\beta H},\quad\text{for }\beta H\gg 1.$ (73)
In slowly driven dynamics, the equilibrium magnetization $m_{\mathrm{eq}}(H(t))$ changes gradually with time. However, in the Curie-Weiss model, we assume that the variation of $m_{\mathrm{eq}}$ is slow enough to neglect its velocity compared to the system’s intrinsic relaxation rate. Therefore, the dominant timescale is set by $\tau_{\mathrm{relax}}$. Let $\tau_{c}$ denote the characteristic time at which the system reaches $m=0$ during the magnetization reversal process. Since $m=0$ is a nonequilibrium value far from the equilibrium $m_{\mathrm{eq}}\approx 1$ under strong field, we estimate this time by
$\tau_{c}\sim e^{-\beta H_{c}},$ (74)
Combining the two expressions gives the scaling relation:
$H_{c}e^{\beta H_{c}}\sim v_{H},$ (75)
with the Lambert W function being the inverse function of $xe^{x}$.
SII.5 $\sigma\sim N^{-1/2}$ correspondence
Substituting Eq. (44) into Eq. (42), the master equation for the magnetization density $m$ in the Curie-Weiss model reads
$\frac{\partial P(m,t)}{\partial t}=$ $\frac{N}{2\tau_{0}}\left[1-\left(m-\frac{2}{N}\right)\right]e^{\beta\left[J(m-2/N+1/N)+H\right]}P(m-\frac{2}{N},t)-\frac{N}{2\tau_{0}}\left[1-m\right]e^{\beta\left[J(m+1/N)+H\right]}P(m,t)$ (76)
$\frac{N}{2\tau_{0}}\left[1+\left(m+\frac{2}{N}\right)\right]e^{-\beta\left[J(m+2/N-1/N)+H\right]}P(m+\frac{2}{N},t)-\frac{N}{2\tau_{0}}\left[1+m\right]e^{-\beta\left[J(m-1/N)+H\right]}P(m,t).$
Using the translation operator, the above equation becomes
$\frac{\partial P(m,t)}{\partial t}=\frac{N}{2\tau_{0}}\left(e^{-\frac{2}{N}\frac{\partial}{\partial m}}-1\right)(1-m)e^{\beta\left[Jm+J/N+H\right]}P(m,t)+\frac{N}{2\tau_{0}}\left(e^{\frac{2}{N}\frac{\partial}{\partial m}}-1\right)(1+m)e^{-\beta\left[Jm-J/N+H\right]}P(m,t).$ (77)
Considering $N\gg 1$, and retaining the $1/N$ term in the above equation, one obtains Eq. (A4) in the end matter of the main text, which returns to the deterministic equation (51) in the $N\to\infty$ limit.
We expand the obtained Fokker-Planck equation [Eq. (A4) in the end matter] around the spinodal point (53) and retain the leading order of each term. For the first drift part in the Fokker-Planck equation, $T(m,H)$ becomes
$T(\hat{m},\hat{H})\approx\beta J\sqrt{\beta J-1}\hat{m}^{2}+\sqrt{\frac{\beta}{J}}\hat{H}.$ (78)
For the second diffusion part in the Fokker-Planck equation, $S(m,H)$ becomes
$S(\hat{m},\hat{H})\approx\frac{1}{\sqrt{\beta J}}.$ (79)
Then we obtain Eq. A(7) in the end matter.
The Fokker-Planck equation for the $\phi^{4}$ model near its spinodal point reads
$\frac{\partial P(\varphi,t)}{\partial t}=-\lambda\frac{\partial}{\partial\varphi}\left[\left(-a_{3}\varphi^{2}+h\right)P(\varphi,t)\right]+\lambda\sigma^{2}\frac{\partial^{2}}{\partial\varphi^{2}}P(\varphi,t),$ (80)
with $\varphi$, $h$ and $a_{3}$ defined in the main text. Comparing the coefficient of the second diffusion part in the above equation and that in Eq. A(7) in the end matter, one obtains $\sigma^{2}\sim N^{-1}$.